# A Compact Post-quantum Strong Designated Verifier Signature Scheme from Isogenies

### v1

@@ 4 Efficiency Analysis
In this section, we analyze the communication and storage overhead of the proposed CSI-SDVS scheme.
The secret keys $\mathsf{sk}_{\mathsf{S}}=\{s_{i}\bmod N\}_{i=1}^{\eta}$ and $\mathsf{sk}_{\mathsf{V}}=\{v_{i}\bmod N\}_{i=1}^{\eta}$, held by the signer and the designated verifier respectively, consist of $\eta$ elements in $\mathbb{Z}_{N}$. Since $\log(N)=\lambda$, the total size of each secret key is $\eta\cdot\lambda=\mathcal{O}(\lambda)$ bits.
Each curve in $\mathcal{E}ll_{p}(\mathfrak{O})$ is uniquely represented by a Montgomery coefficient $A\in\mathbb{F}_{p}$, defining a curve $E_{A}:y^{2}=x^{3}+Ax^{2}+x$. Hence, the public keys $\mathsf{pk}_{\mathsf{S}}$ and $\mathsf{pk}_{\mathsf{V}}$ each contain $\eta$ such elements over $\mathbb{F}_{p}$, yielding a size of $\eta\log(p)=\mathcal{O}(\lambda)$ bits.
A signature $\sigma=(h,\mathbf{z})$, where $\mathbf{z}=\{z_{i}\bmod N\}_{i=1}^{\eta}$, consists of a $\lambda$-bit hash and $\eta$ elements in $\mathbb{Z}_{N}$, each of $\lambda$ bits. This results in a total signature size of $\lambda(1+\eta)=\mathcal{O}(\lambda)$ bits.
Following CSI‑FiSh and CSI‑SharK [33], we adopt $\eta=16$ for 128-bit security, striking a balance between soundness error, key size, and efficiency. Higher values such as $\eta=32$ may be chosen to further reduce the soundness error at the expense of increased overhead.
In comparison, lattice-based SDVS schemes [14, 15, 16, 17] typically require key and signature sizes of $\mathcal{O}(\lambda^{2})$ due to high-dimensional trapdoor structures. Our isogeny-based construction avoids such overhead by relying on group actions in fixed-size cyclic groups, achieving compact $\mathcal{O}(\lambda)$ complexity across all components.
@@ 5 Security Analysis of CSI-SDVS
We prove that the CSI-SDVS scheme satisfies SUF-CMA security under Definition 2.2. Suppose, for the sake of contradiction, that there exists a PPT adversary $\mathcal{A}$ that succeeds in the experiment $\mathsf{Exp}^{\mathsf{SUF\text{-}CMA}}_{\mathsf{CSI\text{-}SDVS},\mathcal{A}}(\lambda)$ (illustrated in Figure 1) with non-negligible probability. Acting as the challenger, a simulator $\mathcal{S}$ interacts with $\mathcal{A}$ by responding to its queries through the signing oracle $\mathcal{O}_{\mathsf{Sign}}$, the simulating oracle $\mathcal{O}_{\mathsf{Simul}}$, the verification oracle $\mathcal{O}_{\mathsf{Verify}}$, and the random oracle $\mathcal{O}_{\mathcal{H}}$. At the end of the interaction, $\mathcal{A}$ outputs a forged signature $\sigma^{\ast}:=(h^{\ast},\mathbf{z}^{\ast})$ on a message $m^{\ast}$ with non-negligible probability, such that
$Y^{\ast}_{i}=[v_{i}+z^{\ast}_{i}]E_{i},\quad\text{and}\quad h^{\ast}=\mathcal{H}(Y^{\ast}_{1}\|\dots\|Y^{\ast}_{\eta}\|m^{\ast})\quad\text{for all }i\in[\eta].$
Leveraging the forgery $\sigma^{\ast}$, the simulator $\mathcal{S}$ computes the secret curves $[s_{i}+v_{i}]E_{0}$ from the public values $[s_{i}]E_{0}$ and $[v_{i}]E_{0}$. This contradicts the assumed hardness of the parallelization problem in the HHS model, as defined in Section 2.4.
@@
    • Extract $\{\bar{E}_{i}\}_{i=1}^{\eta}$ from $(m,\{Y_{i}\}_{i=1}^{\eta},h,\{\bar{E}_{i}\}_{i=1}^{\eta})$, and check if $[z_{i}]E_{i}=\bar{E}_{i}$ or for all $i\in[\eta]$.
@@
Extracting the Forgery. Eventually, the adversary $\mathcal{A}$ outputs a forgery $(m^{\ast},\sigma^{\ast})$ with non-negligible probability, where $\sigma^{\ast}:=(h^{\ast},\mathbf{z}^{\ast})$. The simulator $\mathcal{S}$ then searches for a matching entry $(m,h,\{Y_{i}\}_{i=1}^{\eta},\perp)$ in the hash list $\mathsf{HList}$ such that $m^{\ast}=m$, $h^{\ast}=h$. If such an entry is found, $\mathcal{S}$ can solve an instance of the parallelization problem in the HHS model, as given in Section 2.4, with a non-negligible probability by recovering the secret value $[s_{i}+v_{i}]E_{0}$ as follows:
$[-z^{\ast}_{i}]Y_{i}=[s_{i}-b^{\ast}_{i}][b^{\ast}_{i}]\hat{E}_{i}=[s_{i}]\hat{E}_{i}=[s_{i}+v_{i}]E_{0},\quad\text{for each }i\in[\eta],$
where $\{s_{i}\}_{i=1}^{\eta}$ and $\{v_{i}\}_{i=1}^{\eta}$ are the secret keys of the signer and verifier, respectively. ∎
@@
    $[z^{(0)}_{i}]E_{i}=[b^{(0)}_{i}-s_{i}]E_{i}=[b^{(0)}_{i}]E_{0}\quad\text{and}\quad[z^{(1)}_{i}]E_{i}=[b^{(1)}_{i}-v_{i}]E_{i}=[b^{(1)}_{i}-v_{i}+s_{i}]E_{0},$
@@
    are uniformly distributed over $\mathbb{Z}_{N}$. Therefore, $[b^{(0)}_{i}]E_{0}$ and $[b^{(1)}_{i}-v_{i}+s_{i}]E_{0}$ are both uniformly distributed over $\mathcal{E}ll_{p}(\mathfrak{O})$.
@@
  • Verification Oracle $\mathcal{O}_{\mathsf{Verify}}$: Upon receiving queries on message-signature pair $(m,\sigma)$ and bit $b\in\{0,1\}$ from $\mathcal{A}$, $\mathcal{C}$ runs: $\mathsf{Verify}(\mathsf{pp},\mathsf{sk_{V}},\mathsf{pk_{S_{b}}},\sigma,m)$ and returns $1$ if the signature is valid and 00 otherwise.
@@
in Step 3. As a result, the adversary cannot determine which signer produced the signature, as doing so would require recovering the curves $Y^{(b)}_{i}$ in order to compute $[s^{(b)}_{i}+v_{i}]E_{0}$:
@@ Conclusion
In this work, we have introduced CSI-SDVS, a novel post-quantum Strong Designated Verifier Signature (SDVS) scheme founded on the class group action over the $\mathbb{F}_{p}$-isomorphism classes of supersingular elliptic curves, inspired by CSIDH and CSI-FiSh. Our construction achieves a compelling balance between security and efficiency, offering strong cryptographic guarantees such as Strong Unforgeability under Chosen-Message Attacks (SUF-CMA), Non-Transferability (NT), and Privacy of Signer’s Identity (PSI). One of the notable advantages of CSI-SDVS lies in its asymptotic efficiency: it achieves key and signature sizes of complexity $\mathcal{O}(\lambda)$, marking a significant improvement over the conventional $\mathcal{O}(\lambda^{2})$ bounds typical of lattice-based SDVS constructions. Although an earlier isogeny-based SDVS scheme [18] was proposed, it has been rendered insecure due to devastating cryptanalytic attacks [20, 21, 22] on SIDH. To the best of our knowledge, CSI-SDVS is currently the only isogeny-based SDVS construction that remains secure against all known quantum and classical attacks while offering both compactness and efficiency.

### v3

@@ 4 Efficiency Analysis
In this section, we analyze the communication and storage overhead of CSI-SDVS. The parameter $\eta\geq 1$ represents the number of parallel instances in the base construction. While $\eta=1$ is considered sufficient for current parameter choices, keeping $\eta$ explicit provides several advantages. These include flexibility for potential future instantiations with $\eta>1$, stronger resistance against multi-target attacks due to increased hash input entropy, and a unified framework for security analysis.
For a security level of $\lambda=128$, we have the following:
The secret keys $\mathsf{sk}_{\mathsf{S}}=\{s_{i}\bmod N\}_{i=1}^{\eta}$ and $\mathsf{sk}_{\mathsf{V}}=\{v_{i}\bmod N\}_{i=1}^{\eta}$, held by the signer and designated verifier respectively, consist of $\eta$ elements in $\mathbb{Z}_{N}$. Since $\log_{2}N=2\lambda$, then the total size of each secret key is $\eta\log_{2}N=\mathcal{O}(\lambda)$ bits.
Each curve in $\mathcal{E}ll_{p}(\mathfrak{O})$ is uniquely represented by a Montgomery coefficient $A\in\mathbb{F}_{p}$, defining a curve $E_{A}:y^{2}=x^{3}+Ax^{2}+x$. Consequently, the public keys $\mathsf{pk}_{\mathsf{S}}$ and $\mathsf{pk}_{\mathsf{V}}$ each contains $\eta$ such elements over $\mathbb{F}_{p}$, yielding a size of $\eta\log_{2}p=\mathcal{O}(\lambda)$ bits.
A signature $\sigma=(h,\mathbf{z})$, where $\mathbf{z}=\{z_{i}\bmod N\}_{i=1}^{\eta}$, consists of a $\lambda$-bit hash together with $\eta$ elements in $\mathbb{Z}_{N}$, each requiring at most $2\lambda$ bits. Hence, the total signature size is $(2\eta+1)\lambda=\mathcal{O}(\lambda)$ bits.
@@ 5 Security Analysis of CSI-SDVS
We prove that the CSI-SDVS scheme satisfies SUF-CMA security under Definition 2.2. Suppose, for the sake of contradiction, that there exists a PPT adversary $\mathcal{A}$ that succeeds in the experiment $\mathsf{Exp}^{\mathsf{SUF\text{-}CMA}}_{\mathsf{CSI\text{-}SDVS},\mathcal{A}}(\lambda)$ (illustrated in Figure 1) with non-negligible probability. Strictly speaking, the simulator $\mathcal{S}$ first interacts with its challenger $\mathcal{C}$ to obtain the public keys of the signer and the designated verifier, and subsequently simulates all oracle responses for $\mathcal{A}$ without ever accessing the corresponding secret keys. For brevity and without loss of generality, we present the proof as if $\mathcal{S}$ itself plays the role of the challenger, directly answering $\mathcal{A}$’s queries via the signing oracle $\mathcal{O}_{\mathsf{Sign}}$, the simulating oracle $\mathcal{O}_{\mathsf{Simul}}$, the verification oracle $\mathcal{O}_{\mathsf{Verify}}$, and the random oracle $\mathcal{O}_{\mathcal{H}}$.
At the end of the interaction, $\mathcal{A}$ outputs a forged signature $\sigma^{\ast}:=(h^{\ast},\mathbf{z}^{\ast})$ on a message $m^{\ast}$ with non-negligible probability, satisfying
$\forall i\in[\eta]\quad Y^{\ast}_{i}=[v_{i}+z^{\ast}_{i}]E_{i},\quad\text{and}\quad h^{\ast}=\mathcal{H}(Y^{\ast}_{1}\|\dots\|Y^{\ast}_{\eta}\|m^{\ast}).$
Leveraging the forgery $\sigma^{\ast}$, the simulator $\mathcal{S}$ computes the secret curves $s_{i}+v_{i}]E_{0}$ from the public values $s_{i}]E_{0}$ and $v_{i}]E_{0}$, thereby breaking the assumed hardness of the parallelization problem in the HHS model, as defined in Section 2.4.
@@
    • Extract $\{\bar{E}_{i}\}_{i=1}^{\eta}$ from $(m,\{Y_{i}\}_{i=1}^{\eta},h,\{\bar{E}_{i}\}_{i=1}^{\eta})$, and check if $z_{i}]E_{i}=\bar{E}_{i}$ or for all $i\in[\eta]$.
@@
Extracting the Forgery. Eventually, the adversary $\mathcal{A}$ outputs a forgery $(m^{\ast},\sigma^{\ast})$ with non-negligible probability, where $\sigma^{\ast}:=(h^{\ast},\mathbf{z}^{\ast})$. The simulator $\mathcal{S}$ then searches for a matching entry $(m,h,\{Y_{i}\}_{i=1}^{\eta},\perp)$ in the hash list $\mathsf{HList}$ such that $m^{\ast}=m$, $h^{\ast}=h$, and the last component is $\perp$. If such an entry is found, $\mathcal{S}$ can solve an instance of the parallelization problem in the HHS model, as given in Section 2.4, with a non-negligible probability by recovering the secret value $s_{i}+v_{i}]E_{0}$ as follows:
$z^{\ast}_{i}]Y_{i}=[s_{i}-b^{\ast}_{i}][b^{\ast}_{i}]\hat{E}_{i}=[s_{i}]\hat{E}_{i}=[s_{i}+v_{i}]E_{0},\quad\text{for each }i\in[\eta],$
which is induced by the secret keys $\{s_{i}\}_{i=1}^{\eta}$ and $\{v_{i}\}_{i=1}^{\eta}$ of the signer and designated verifier, respectively. ∎
@@
    $z^{(0)}_{i}]E_{i}=[b^{(0)}_{i}-s_{i}]E_{i}=[b^{(0)}_{i}]E_{0}\quad\text{and}\quad[z^{(1)}_{i}]E_{i}=[b^{(1)}_{i}-v_{i}]E_{i}=[b^{(1)}_{i}-v_{i}+s_{i}]E_{0},$
@@
    are uniformly distributed over $\mathbb{Z}_{N}$. Therefore, $b^{(0)}_{i}]E_{0}$ and $b^{(1)}_{i}-v_{i}+s_{i}]E_{0}$ are both uniformly distributed over $\mathcal{E}ll_{p}(\mathfrak{O})$.
@@
  • Verification Oracle $\mathcal{O}_{\mathsf{Verify}}$: Upon receiving queries on message-signature pair $(m,\sigma)$ and bit $b\in\{0,1\}$ from $\mathcal{A}$, $\mathcal{C}$ runs: $\mathsf{Verify}(\mathsf{pp},\mathsf{sk_{V}},\mathsf{pk_{S_{b}}},\sigma,m)$ and returns $1$ if the signature is valid and 0 otherwise.
@@
in Step 3. As a result, the adversary cannot determine which signer produced the signature, as doing so would require recovering the curves $Y^{(b)}_{i}$ in order to compute $s^{(b)}_{i}+v_{i}]E_{0}$:
@@ Conclusion
In this work, we introduced $\mathsf{CSI\text{-}SDVS}$, a new post-quantum Strong Designated Verifier Signature (SDVS) scheme based on isogenies. By leveraging the class group action on supersingular elliptic curves and the hardness of the Multi-Target Group Action Inverse Problem (MT-GAIP), our construction achieves Strong Unforgeability (SUF-CMA), Non-Transferability (NT), and Privacy of the Signer’s Identity (PSI) in the random oracle model. Compared to lattice-based SDVS schemes [14, 15, 16, 17], which typically require larger key and signature sizes due to high-dimensional trapdoor structures, $\mathsf{CSI\text{-}SDVS}$ achieves compact key and signature sizes, making it practical for resource-constrained environments. Although an earlier isogeny-based SDVS scheme [18] was proposed, it has been rendered insecure due to devastating cryptanalytic attacks [20, 21, 22] on SIDH. To the best of our knowledge, CSI-SDVS is the only isogeny-based SDVS scheme that achieves security against known quantum and classical attacks, along with compactness and efficiency.